At school, there are three bells. The first bell rings every 20 minutes, the second rings every 30 minutes, and the third rings every 45 minutes. All three bells rang at exactly 10 a.m. At what other times will the three bells will ring at the same time? Select all that apply.

Answer by stanbon(75887)

At school, there are three bells. The first bell rings every 20 minutes, the second rings every 30 minutes, and the third rings every 45 minutes. All three bells rang at exactly 10 a.m. At what other times will the three bells will ring at the same time? Select all that apply.
----------------
Find the least common multiple of 20, 30, and 45
-----
Prime-factor Form::
20 = 2^2 * 5
30 = 2*3 * 5
45 = 3^2 * 5
------
Least common multiple:: 2^3*3^2*5 = 8*9*5 = 360 minutes
------
Note: 360 minutes = 6 hrs
-----
Ans: 4pm , 10pm, 4am, 10am
================================
Cheers,
Stan H.
----------------